you sell 500 bags of popcorn per day at the student union. the cost of the kernels is $.08 per bag; the paper bag costs $.04, you estimate the cost of cooking oil and salt per bag is $.02, and rental for the space to put the machine is $50/day. If you want to make a 80% gross margin on popcorn, what do you need to price it at per bag?
Answer #1

rental space expense per bag = $50 / 500 = $0.1

Total cost of one one popcorn = cost of the kernels + the paper bag cost + cost of cooking oil and salt per bag + rental space expense per bag = 0.08 + 0.04 + 0.02 + 0.10 = $0.24

If you want 80% gross margin then price of popcorn per bag = $0.24 / 1 - 0.80 = $0.24 / 0.20 = $1.20